OpenStudy (anonymous):

Which equations in point-slope form are equations of the line that pass through the points (9, −2) and (−3, 6)? Choose exactly two answers that are correct. A. B. C. D.

OpenStudy (anonymous):

a. y + 2 = - 2/3 (x-9) b. y - 9 = - 3/2 (x + 2) c. y + 1 = - 3/2 (x - 4) d. y - 6 = - 2/3 (x + 3)

OpenStudy (anonymous):

hhhhhhhhhhhhhhhiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiii :D

OpenStudy (anonymous):

um... do you need help?

OpenStudy (anonymous):

eehh what the hell ill answer it anyway :D so we begin by finding the slope... y2-y1/x2-x1 so we plug in our points, (9,-2) (-3,6) (x1,y1) (x2,y2)

OpenStudy (anonymous):

6-(-2)/-3-9 8/-12

OpenStudy (anonymous):

sorry i wasnt in the room but thank you and heres your medal

OpenStudy (anonymous):

wait hold on im confused @atreyu6s3x6 which to is it then

OpenStudy (anonymous):

two not to stupid auto correct

OpenStudy (anonymous):

so if we reduce it we get -2/3 as our slope :)

OpenStudy (anonymous):

so a and d ?

OpenStudy (anonymous):

yaaasss!! :D

OpenStudy (anonymous):

thank you i really needed that.

OpenStudy (anonymous):

no problem :D have a great day